The current value of a vehicle is ksh.128,000.If this vehicle has been depreciating by a fifth of its value every year, find its value 4 years ago.
Answer by Boreal(15235)

original was x
x*(1-0.2)^4, is depreciation of 20% or a fifth, for 4 years
x*0.8^4=128000
.4096x=128000
x=128000/0.4096=312,500
if one multiplies 312,500*0.8, and the answer by 0.8, and that answer by 0.8, and that answer by 0.8=128000